OpenStudy (anonymous):

laplace((1/s-e^(-2s))/(s^2+2s+2)

OpenStudy (anonymous):

plis help me...

OpenStudy (anonymous):

thanks....really help me...

OpenStudy (anonymous):

(e^x cos 6x)(6e^x cos6x)

OpenStudy (anonymous):

help me..multiplication of it?

OpenStudy (anonymous):

it would be \[6e^{2x}\cos ^{2}(6x)\]

OpenStudy (anonymous):

this one (e^x cos 6x)(e^x sin 6x)

OpenStudy (turingtest):

You want the Laplace transform of\[e^{2x}\sin(6x)\cos(6x)\]??

OpenStudy (anonymous):

nope just the multiplication of that?

OpenStudy (anonymous):

owh that the answers ya..

OpenStudy (anonymous):

turing test already did the multiplication of it

OpenStudy (anonymous):

owh ok2..i got it..thanks...

OpenStudy (anonymous):

how about this multiplication (e^x/3cos x)(6e^xcos 6x+e^xsin 6x)???